Consider a simple model where workers can be either employed or unemployed. If the job separation rate is 2% per month and the job finding rate is 53% per month, which of the following is FALSE?

a) There are more workers losing their jobs than there are unemployed workers finding a job

b) None of the other options

c) The steady state unemployment rate is 3.6%

d)The steady state employment rate is 96.4%
Answer #1

Job seperation rate, is when people lose their jobs , either voluntarly or involuntarily.

Job finding rate is the rate of unemployed people finding jobs.

Now, as there can be only two kinds of workers, either employed or unemployed.

So, its obvious and logical that the people who get jobs, replace those who lose their jobs. So, the the number of workers who are losing their jobs is equal to the number of unemployed workers who are finding a job.

So, the statement a) is FALSE.

The steady state unemployed rate can be found by the theory that at steady state:

Number of people who lose their jobs = s * E = f * U = number of unemployed people who find jobs

so, U/L = s/(s+f) , where L = E+U, s = job seperation rate, f = job finding rate

steady state unemployment rate = s/(s+f) = 0.02/(0.02+0.53) = 0.02/0.55 * 100 = 3.63%

so, statement c) is TRUE.

steady state employment rate = 100-3.63 =96.37%, as the model consists of only employed and unemployed.

So, stament d) is TRUE.

If statement b) would be true, i.e. it would imply that none of the other options are false. but, we see that statement a) is false. so, none of the other options is false, is also false. So, the statements a) and b) are FALSE.